# Mesoscopic Modeling of High-Density Carbon Nanotube Films for Memristive Device Applications

### v1

@@ Abstract
Carbon nanotube (CNTs) materials, which exhibit intrinsically high electrical conductivity, are promising candidates for energy-efficient electronic devices. Recently, high-density CNT films have also been successfully employed as switching elements in non-volatile memory cells. However, the mechanism of electrical conduction through such complex systems is still poorly understood. To identify structural parameters that govern the electrical current in CNT films, we employed coarse-grained molecular dynamics to construct dense mesoscale CNT film models, where we considered CNTs with different chiralities and lengths. The effects of CNT geometrical features on the film morphologies were quantified by devising a set of structural descriptors and analyzing their mutual correlations. The impact of varying the concentration of amorphous carbon (aC) inclusions on the film structure was assessed. Finally, we employed a nodal analysis framework to compute the electrical current across the networks and correlate the charge transport characteristics to the underlying structural descriptors. Transport is found to be enhanced in films that exhibit high curvature and buckling, low bundling, and strong connectivity, with amorphous carbon components playing a nontrivial configuration-dependent role. These findings provide a framework for the rational design of CNT-based memristor architectures and highlight the potential of mesoscale modeling to guide the engineering of advanced nanostructured materials.
@@
High-Density Carbon Nanotube Films, Memristive Devices, Amorphous Carbon, Mesoscopic Modeling, Electrical Current, Structural Descriptors, Morphology-Transport Correlation1 Introduction
@@
Carbon-based electronics is widely regarded as a promising alternative to conventional silicon-based devices 1, 2, 3, 4, 5, 6, 7. In particular, carbon materials such as carbon nanotubes (CNTs) and graphene exhibit exceptional electronic, optoelectronic and mechanical properties 8, 9, 10. The extremely high predicted tensile strength and Young’s modulus of CNTs make them among the strongest and stiffest known materials 11. Although the properties of single and small ensembles of CNTs have been extensively studied in recent decades, both experimentally and theoretically, the properties of CNT films, which are strongly affected by the formation of bundles and junctions 2, 12, 13, 14, 15, 16, remains comparatively less understood 17, 18.
@@
Experimentally, it has been shown that the properties of CNT films can strongly depend on their density 19, 20, 21, 22, 23, 24. Various strategies have been developed to increase the density of CNT films, such as annealing to increase ordering and crystallinity 19, 21 or mechanical pressure to promote tighter packing 23. For instance, buckypapers (i.e. macroscopic aggregates of carbon nanotubes) exhibit almost linearly improved electrical and thermal conductivity under compression 23. The fabrication method also seems important, as sequential layering of CNT films can be used to tune the properties of the final film 25, 9. Finally, although less frequently discussed, amorphous carbon is expected to be inherently present between tubes and bundles as a byproduct of the CNT growth process, but its effect on the physical properties of CNT films remains poorly understood.
@@
Recently, CNT films have found new applications in resistive random-access memory (RRAM) devices, which are considered promising candidates for the next generation of nonvolatile memory (NRAM) technologies 26, 27. In such devices, the memory state information is encoded in the conductivity of the NRAM cell determined by reversible changes in cell resistance. Resistive switching between these states is governed by a complex interplay between local structural features and electron transport pathways at the nanoscale. Understanding the mechanisms of electron transport through CNT films remains challenging due to their intrinsically disordered structure, heterogeneous connectivity, and possible presence of amorphous carbon.
@@
Mesoscale modeling offers an attractive framework to address these issues 28, 29, 30, 31. It allows the direct incorporation of structural complexity, such as bundling, curvature, and intertube junctions into current-carrying networks, without relying on computationally expensive atomistic simulations or overly simplified continuum models. In particular, the use of structural descriptors may provide a compact and interpretable representation of key morphological features, thereby enabling a systematic exploration of how geometry and topology influence electron transport.
@@
Early theoretical work has highlighted the importance of geometric morphology in determining the physical properties of CNT films 32, 33. However, these studies focused on two-dimensional (2D) networks and did not include an atomistic description of the film structure. More recently, computational modeling has been used to understand the behavior of CNT films at a mesoscopic level. The methodology can generally be divided into two classes: (i) the use of a mesoscopic tubular (MT) potential that describes inter-CNT interactions based on a continuous interatomic potential 28, 34, 30, 35, and (ii) the mesoscopic distinct element method (MDEM), which enables the simulation of interactions between particles of arbitrary shapes 36, 37, 38, 39. Simulations using these mesoscopic models have shown that various factors, particularly those related to underlying structure of the network, have a strong influence on mechanical and thermal properties 34, 40, 31, 35, 39. In contrast to the 2D approximation representing CNTs as straight lines, the MT potential allows one to directly study the effects of CNT bending and bundling, as well as the construction of more complex structures 29, 30, 34, 41, 42, 43. For example, it has been shown that increased bending energy leads to an overall increase in film stability, defined as the tendency not to disintegrate into cellular structures consisting of individual bundles weakly connected to each other 34.
@@
In this study, we construct and analyze high-density CNT films with varying chirality, length, density, and aC content to investigate whether such films can be rationally designed to achieve specific target morphologies. Unlike previous studies that focused on relatively low-density systems, we consider films with density approaching the close-packed limit and study both single-layer and stacked multi-layer configurations, thereby allowing us to assess the influence of the fabrication method. We employ a mesoscale modeling approach capable of capturing mechanical deformation and local disorder. By quantifying each structure through a set of physically motivated descriptors, we establish correlations between structural properties and electron transport efficiency. Moreover, we develop a simple approach to mimic the effect of amorphous carbon inside the films. Overall, this work sheds light on the structural factors governing conductivity in complex CNT networks and provides a foundation for future modeling of resistive switching mechanisms in CNT-based memristive devices.
@@
The paper is organized as follows. We discuss the choice of structural parameters, such as tube length and film density, in Section 2.1. The mesoscopic model and the incorporation of aC into the simulations are presented in Section 2.2. The procedure for constructing the films is presented in Section 2.3. We define the set of descriptors in Section 2.4. The nodal analysis used to compute the electrical current through the films is detailed in Section 2.5. The results of the simulations are presented in Section 3 followed by concluding remarks in Section 4. Additional data and supporting figures are provided in the Supplemental Information (SI).
@@ 2.1 Structural parameters
The CNT films targeted in this study are modeled after experimental systems fabricated as described in Refs 27, 44. Due to the specifics of the growth and deposition techniques, the formed CNT films consist of stacked layers with different CNT lengths and chirality distributions (see SI). Based on the experimentally measured distributions, we selected three representative lengths: a short one (15 nm), a intermediate one (40 nm), and a long one (100 nm, corresponding to half the size of the simulation box), as well as two different chiralities, (16,0) and (32,0), each associated with a specific diameter: and .
Experimentally fabricated films are predominantly composed of multi-walled CNTs, whereas simulated films are composed of single-walled CNTs. We chose a high density of 0.6 gcm-3 (the close-packed density for (32,0) tubes is 0.8 gcm-3), and also simulate films at the lower density of 0.3 gcm-3 for comparison. The films were constructed either as single-layer or as four-layer stacked configurations. Finally, the amorphous carbon content was experimentally estimated to lie between 10 and 40% by mass. Starting from two reference configurations, namely 100 nm CNT length, single-layer, 0.6 gcm-3, and either (16,0) or (32,0), the amount of aC is varied from 0% to 40% with a 10% increment. By keeping constant the target density, the number of CNTs in the simulation box decreases as the amount of aC particles increases. The standard deviation of the final density is 710-4 gcm-3 for the 0.3 gcm-3 samples and 410-4 gcm-3 for the 0.6 gcm-3 samples. Altogether, a total of 32 distinct structures were constructed and analyzed.
@@ 2.2 Mesoscopic Model
CNTs are modeled as collections of point particles (beads), each representing a 1.0 nm-long cylindrical segment of a CNT; the terminal beads are lighter and only represent a 0.5-nm segment. Beads within the same CNT are connected by harmonic stretching and bending potentials, while beads from different, randomly oriented CNTs interact through nonbonded Lennard-Jones potentials, according to the mesoscopic tubular force field developed by Volkov, Zhigilei, and collaborators 28, 34, 30, 29. Only single-walled CNTs are considered in this work. All calculations are performed with the MESONT package 45, 46 as implemented in the LAMMPS code 47.
@@
Amorphous carbon particles are also represented as single beads interacting with the other entities in the system (CNTs and other aC particles) via nonbonded Lennard-Jones potentials. Their interactions are parametrized with the same procedure used for CNTs, starting from the airebo potential 48. To this end, aC particles are first modeled as a collection of carbon atoms, randomly distributed within a spherical volume, with a minimum C–C distance corresponding to a double bond (0.134 nm) (see Fig. 1). The sphere radius is set to , and the number of carbon atoms is adjusted to achieve a 2.5 gcm-3 density. The particle is then placed at varying distances from another entity, and the total interaction energy is computed as the sum of all pairwise C–C airebo contributions. This calculation is repeated 300 times with different random internal configurations of the aC structure. The resulting interaction energy profiles are averaged and fitted to determine the two Lennard-Jones parameters and used in the classical force field (see Table 1).
@@
In this work, a single particle size and fixed density are adopted for all aC beads. Nonetheless, the same parametrization scheme can be extended to generate interaction parameters for a broad range of particle sizes and internal densities, thereby enabling flexible modeling of diverse aC morphologies relevant to different experimental conditions and device architectures.
2.3 CNT films
@@
Film models are generated by initially placing straight CNTs and aC particles at random positions and orientations within a large simulation box, corresponding to a low-density initial configuration. The dimensions and remain fixed throughout the procedure, while only the dimension is varied during compression. The system is confined between two parallel repulsive walls in the direction, whose interactions with the particles are described by a 12-6 Lennard-Jones potentials. These walls represent the physical confinement imposed by the top and bottom electrodes. Periodic boundary conditions (PBCs) are applied only in the -plane.
@@
To reach the target film density of either 0.3 or 0.6 gcm-3, vertical downward forces are applied to all entities in the initially expanded box while gradually moving downward the upper wall (see SI for a description of the compression protocol). The dynamics of all particles is integrated in the canonical () ensemble, with the temperature controlled by a Nosé-Hoover thermostat set at . Once the desired volume is reached, the system is equilibrated at constant temperature until the total energy fluctuates around a steady negative value, after which a final energy minimization is performed to reach a local minimum. This procedure produces a single-layer CNT film contained within a simulation box.
@@
Multi-layer films are generated by first equilibrating four low-density layers which are then stacked together before the compression protocol (see SI), thereby yielding films with the identical overall dimensions and density that their single-layer counterparts. As discussed in section 3, structures with higher density and CNT diameter are more difficult to compact. For these cases, several compression-decompression cycles with intermediate annealing stages at elevated temperature were required to reach the desired volume.
@@ 2.4 Structural metrics
We used structural descriptors or metrics that offer a compact and interpretable representation of key morphological features and enable systematic exploration of how geometry and topology influence electron transport. The first metric quantifies the orientation of the CNTs with respect to the -plane. Because CNTs may bend and buckle, the analysis is performed on each of their segments separately. From the unitary direction vector of a segment, the angle with the -plane is calculated from its component:
@@
where is the number of buckling points within the -th CNT containing beads. is the total number of CNTs in the film. CNT beads are identified as buckling points when the segments they connect are misaligned by more than 28.
The third metric is the curvature factor , which quantifies the average curvature of CNTs (excluding sharp kinks). First we calculate the geometric curvature of the -th CNT, at the position of its -th bead:
where the derivative and are obtained using the finite differences method:
@@
where is the total number of segments, the number of connected components, and the size (number of segments) of the -th component. This metric reflects the average number of connected neighbors per segment and captures local and global bundling in the film. We also report the average bundle size as an additional indicator of aggregate structure. In the limiting case where all CNT segments belong to a single connected component (i.e.: one large bundle spanning the entire system) and assuming that PBCs also apply in the -direction, the bundling metric would reach its maximum value . In the opposite case where no segment is connected to any other (i.e.: all segments are isolated) one obtains .
@@
where is the number of internal junctions in the path. This descriptor favors short and direct conduction pathways. Second, minimum number of junctions corresponds to the smallest number of internal junctions observed across all conductive paths and provides a lower bound on the topological complexity required for charge transport through the network. These descriptors allow us to distinguish between highly interconnected films supporting efficient conduction and sparse networks where charge must percolate through long paths. The last descriptor is simply the smallest number of connections with either the bottom or the top electrode.

### v2

# Mesoscopic Modeling of Structure-Transport Relationships in Dense CNT Films Containing Amorphous Carbon

@@ Abstract
Carbon nanotube (CNT) films are widely considered as prospective building blocks for advanced electronic and nanostructured materials. In particular, electrical transport in high-density CNT films results from a complex interplay between network morphology and CNT connectivity, which remains challenging to characterize quantitatively. To identify the structural parameters that govern the electrical current in CNT films, we employed coarse-grained molecular dynamics to construct dense mesoscale CNT film models that include CNTs with different chiralities and lengths. The effects of CNT geometrical features on the film morphologies were quantified by devising a set of structural descriptors and analyzing their mutual correlations. The impact of varying the concentration of amorphous carbon (aC) inclusions on the film structure was assessed. Finally, we employed a nodal analysis framework to compute the electrical current across the networks and correlate the charge transport characteristics to the underlying structural descriptors. The current is found to be enhanced in films that exhibit high curvature and buckling, low bundling, and strong connectivity. We discuss how the presence of aC inclusions modifies these morphological and current characteristics. This work provides a mesoscale modeling framework for modeling structure–transport relationships in dense CNT films and highlights the role of morphological descriptors in guiding the interpretation of electrical transport in complex nanostructured networks.
@@
High-Density Carbon Nanotube Films, Amorphous Carbon, Mesoscopic Modeling, Electrical Current, Structural Descriptors, Morphology-Transport Correlation1 Introduction
@@
Carbon-based electronics is widely regarded as a promising alternative to conventional silicon-based devices 3, 21, 8, 40, 26, 34, 6, 70. In particular, carbon materials such as carbon nanotubes (CNTs) and graphene exhibit exceptional electronic, optoelectronic and mechanical properties 10, 55, 30. The extremely high predicted tensile strength and Young’s modulus of CNTs make them among the strongest and stiffest known materials 47. Although the properties of single CNTs and small ensembles of CNTs have been extensively studied in recent decades, both experimentally and theoretically, the properties of CNT films, which are strongly affected by the formation of bundles and junctions 21, 12, 13, 64, 56, 27, remain comparatively less understood 9, 63.
@@
Experimentally, it has been shown that the properties of CNT films can strongly depend on their density 68, 2, 67, 39, 69, 29. Various strategies have been developed to increase the density of CNT films, such as annealing to increase ordering and crystallinity 68, 67 or mechanical pressure to promote tighter packing 69. For instance, buckypapers (i.e. macroscopic aggregates of carbon nanotubes) exhibit almost linear enhancement of electrical and thermal conductivity under compression 69. The fabrication method also plays an important role, as sequential deposition of CNT layers can be used to tune the properties of the final film 41, 55. Finally, the presence of amorphous carbon between and around CNTs is inherent to the CNT growth process, yet its effect on the physical properties of the films remains poorly understood.
@@
Recently, CNT films have emerged as promising candidates for resistive random access memory (RRAM) and neuromorphic applications. 42, 16, 66. In these applications, CNT films are used as a conductive media, however, our understanding of the mechanisms of electron transport through CNT films remains incomplete due to their intrinsically disordered structure, heterogeneous connectivity, and presence of amorphous carbon.
@@
Mesoscale modeling offers an attractive framework to address these issues 71, 51, 54, 53. It allows the direct incorporation of structural complexity, such as bundling, curvature, and inter-tube junctions into current-carrying networks, without relying on computationally expensive atomistic simulations or overly simplified continuum models. In particular, the use of structural descriptors may provide a compact and interpretable representation of key morphological features, thereby enabling a systematic exploration of how geometry and topology influence electron transport.
@@
Early theoretical work has highlighted the importance of structural morphology in determining the physical properties of CNT films 31, 4. However, these studies focused on two-dimensional (2D) networks and did not include an atomistic description of the film structure. More recently, computational modeling has been used to understand the behavior of CNT films at a mesoscopic level. The methodology can generally be divided into two classes: (i) the use of a mesoscopic tubular (MT) potential that describes inter-CNT interactions based on a continuous interatomic potential 71, 52, 54, 50, and (ii) the mesoscopic distinct element method (MDEM), which enables simulation of interactions between particles of arbitrary shapes 36, 57, 37, 58. Simulations using these mesoscopic models have shown that various factors, particularly those related to the underlying structure of the network, strongly affect its mechanical and thermal properties 52, 32, 53, 50, 58. In contrast to the 2D approximation representing CNTs as straight lines, the MT potential allows one to directly study the effects of CNT bending and bundling, as well as construct more complex structures 51, 54, 52, 61, 59, 60. For example, it has been shown that increased bending energy leads to an overall increase in film stability, defined as the tendency not to disintegrate into cellular structures consisting of individual bundles weakly connected to each other 52.
@@
In this study, we focus on creating and analyzing high-density films with different CNT chirality, length, density, and aC content to investigate whether such films can be rationally designed to achieve specific target morphologies. Unlike previous studies that focused on relatively low-density systems, we consider films with density approaching the close-packed limit and study both single-layer and stacked multi-layer structures, similar to those used in experimental studies 16, 43. We employ a mesoscale modeling approach capable of capturing mechanical deformation and local disorder. By quantifying each structure through a set of physically motivated descriptors, we establish correlations between structural properties and electron transport efficiency. Moreover, we develop a simple approach to mimic the structural effect of amorphous carbon particles inside the films. This work sheds light on the structural factors that govern conductivity in complex CNT networks and provides a foundation for future modeling of electrical transport phenomena in dense CNT films and related nanostructured networks.
@@
The paper is organized as follows. We discuss the choice of structural parameters, such as tube length and film density, in Section 2.1. The mesoscopic model and the incorporation of aC into the simulations are presented in Section 2.2. The procedure for constructing the films is presented in Section 2.3. We define the set of descriptors in Section 2.4. The nodal analysis used to compute the electrical current through the films is detailed in Section 2.5. The results of the simulations are presented and discussed in Section 3 followed by discussion in section 4 and concluding remarks in Section 5. Additional data and supporting figures are provided in the Supplementary Material (SM).
@@ 2.1 Structural parameters
Modeling of 3D CNT assemblies is challenging due to the huge variety of CNT types (single- and multi-wall), chirality, lengths, and the complexity of their interaction and the morphology of the films, which includes CNT bending and bundling 18. Driven by particular applications, it is inevitable to consider specific systems, characterized by a particular film density, morphology, and distribution of CNT lengths and chiralities. In particular, the film morphology is directly determined by the method of deposition and treatment. Here, we are guided by CNT films produced for RRAM devices fabricated as described in Refs 16, 43. However, the methodology and analysis will be useful for future simulations of dense CNT films for other nanoelectronic applications.
Due to the specifics of the growth and deposition techniques, the CNT films considered in our work consist of stacked layers with different CNT lengths and chirality distributions (see SM). Based on the experimentally measured distributions, we selected three representative lengths: short (15 nm), intermediate (40 nm), and long (100 nm, corresponding to half the size of the simulation box), as well as two different chiralities, (16,0) and (32,0), each associated with a specific diameter: and . These chiralities are chosen to represent the average diameter of CNTs observed in RRAM devices (see Fig. S2) independently of the metallic or semiconducting character of the CNTs. As discussed below, the nodal analysis used to compute the electrical current depends on CNT diameter only, and not on the specific chiral indices.
Experimentally fabricated films are often composed of multi-walled CNTs, although RRAM devices have also been successfully realized using single-walled CNTs, indicating that multi-walled structures are not a strict requirement for device operation. In the present study, only single-walled CNTs are considered, as interactions between multi-walled CNTs are not currently implemented in the coarse-grained model. We chose a high density of 0.6 gcm-3 (the close-packed density for (32,0) tubes is 0.8 gcm-3), and also simulate films at the lower density of 0.3 gcm-3 for comparison. The films were constructed in single-layer or four-layer stacked configurations. Finally, the amorphous carbon content was experimentally estimated to be between 10 and 40% by mass. Based on two reference configurations, namely 100 nm CNT length, single-layer, 0.6 gcm-3, and either (16,0) or (32,0) chiralities, the amount of aC is varied from 0% to 40% with a 10% increment. By keeping the target density constant, the number of CNTs in the simulation box decreases as the amount of aC particles increases. The standard deviation of the final density is 710-4 gcm-3 for the 0.3 gcm-3 samples and 410-4 gcm-3 for the 0.6 gcm-3 samples. In total, 32 distinct structures were constructed and analyzed.
@@ 2.2 Mesoscopic Model
CNTs are modeled as collections of point particles (beads), each representing a 1.0 nm-long cylindrical segment of a CNT; the terminal beads are lighter and only represent a 0.5-nm segment. Beads within the same CNT are connected by harmonic stretching and bending potentials, while beads from different, randomly oriented CNTs interact through nonbonded Lennard-Jones potentials, according to the mesoscopic tubular force field developed by Volkov, Zhigilei, and collaborators 71, 52, 54, 51. All calculations were performed with the MESONT package 25, 28 as implemented in the LAMMPS code 49.
@@
Amorphous carbon particles are also represented as single beads that interact with the other entities in the system (CNTs and other aC particles) via nonbonded Lennard-Jones potentials. Their interactions are parametrized with the same procedure used for CNTs, starting from the airebo potential 46. To this end, aC particles are first modeled as a collection of carbon atoms, randomly distributed within a spherical volume, with a minimum C–C distance corresponding to a double bond (0.134 nm) (see Fig. 1). The radius of the sphere is set to , and the number of carbon atoms is adjusted to achieve a 2.5 gcm-3 density. The particle is then placed at varying distances from another entity, and the total interaction energy is computed as the sum of all pairwise C–C airebo contributions. This calculation is repeated 300 times with different random internal configurations of the aC structure. The resulting interaction energy profiles are averaged and fitted to determine the two Lennard-Jones parameters and used in the classical force field (see Table 1).
@@
In this work, a single particle size and fixed density are adopted for all aC beads. Nonetheless, the same parametrization scheme can be extended to generate interaction parameters for a broad range of particle sizes and internal densities, enabling flexible modeling of diverse aC morphologies relevant to different experimental conditions and device architectures.
2.3 Creating CNT film models
@@
CNT film models were created by placing straight CNTs and aC particles at random positions and orientations within a large simulation box, corresponding to a low-density initial configuration. The lateral dimensions of the box and remained fixed throughout the procedure, while the dimension was varied during compression to create a dense film. The system was confined between two parallel repulsive walls in the direction, whose interactions with the particles were described by a 12-6 Lennard-Jones potentials. These walls represent the physical confinement imposed by the top and bottom electrodes. Periodic boundary conditions (PBCs) are applied only in the -plane.
@@
To reach the target film density of either 0.3 or 0.6 gcm-3, vertical downward forces are applied to all entities in the simulation box while gradually moving downward the upper wall (see SI for a description of the compression protocol). The dynamics of all particles is integrated in the canonical () ensemble, with the temperature controlled by a Nosé-Hoover thermostat set at . Once the desired volume is reached, the system is equilibrated at constant temperature until the total energy fluctuates around a steady negative value, after which a final energy minimization is performed to reach a local minimum. This procedure produces a single-layer CNT film contained within a simulation box.
@@
Multi-layer films are generated by first equilibrating four low-density layers which are then stacked together before the compression protocol is applied (see SM), thereby yielding films with the identical overall dimensions and density as their single-layer counterparts. As discussed in section 3, structures with higher density and CNT diameter are more difficult to compact. For these cases, several compression-decompression cycles with intermediate annealing stages at elevated temperature were required to reach the desired volume.
@@ 2.4 Structural metrics
To analyze the structure of obtained films, we used structural descriptors or metrics that offer a compact and interpretable representation of key morphological features and enable systematic exploration of how geometry and topology influence electron transport.
The first metric quantifies the orientation of the CNTs with respect to the -plane. Because CNTs may bend and buckle, the analysis is performed on each of their segments separately. From the unitary direction vector of a segment, the angle with the -plane is calculated from its component:
@@
where is the number of buckling points within the -th CNT containing beads. is the total number of CNTs in the film. CNT beads are identified as buckling points when the segments they connect are misaligned by more than 71.
The third metric is the curvature factor , which quantifies the average curvature of CNTs (excluding sharp kinks). First, we calculate the geometric curvature of the -th CNT, at the position of its -th bead:
where the derivatives and are obtained using the finite difference method:
@@
where is the total number of segments, the number of connected components, and the size (number of segments) of the -th component. This metric reflects the average number of connected neighbors per segment and captures local and global bundling in the film. We also report the average bundle size as an additional indicator of the aggregate structure. In the limiting case where all CNT segments belong to a single connected component (i.e.: one large bundle spanning the entire system) and assuming that PBCs also apply in the -direction, the bundling metric would reach its maximum value . In the opposite case where no segment is connected to any other (i.e.: all segments are isolated) one obtains .
@@
where is the number of internal junctions in the path. This descriptor favors short and direct conduction pathways. Second, the minimum number of junctions corresponds to the smallest number of internal junctions observed across all conductive paths and provides a lower bound on the topological complexity required for charge transport through the network. These descriptors allow us to distinguish between highly interconnected films supporting efficient conduction and sparse networks, where the charge must percolate through long paths. The last descriptor is simply the smallest number of connections with the bottom or the top electrode.